### v2

@@ 1 Purpose and outline
Since its proposal, the fuzzy sphere regularisation has made significant contributions to the study of 3d CFTs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The Julia package FuzzifiED aims at simplifying the numerical calculations on the fuzzy sphere. It supports exact diagonalisation (ED) calculations, as well as the density matrix renormalisation group (DMRG) using the ITensor [20] library. FuzzifiED can also apply to generic fermionic and bosonic models. This package offers the following features :
@@
    Versatility : FuzzifiED can reproduce nearly all the ED and DMRG results from fuzzy sphere research. Its flexible design also makes it straightforward to adapt to new models.
@@
  • In Section 4, we review the setup of fuzzy sphere, the construction of interaction models and the extraction of CFT data.
@@
  • In Appendix B, we present the construction of the Landau level and pseudopotentials on the torus.
@@
Part I Review of fuzzy sphere
@@ 2.1 Conformal field theory
Conformal field theory (CFT) is one of the central topics of modern physics. It refers to a field theory that is invariant under conformal transformations that preserve the angles between vectors. In spacetime dimension $d>2$, the global conformal symmetry transformations form a group $\mathrm{SO}(d+1,1)$, generated by translation, $\mathrm{SO}(d)$ rotation [^1] Here we work in Euclidean signature. In Lorentzian signature it is the Lorentz transformation $\mathrm{SO}(1,d-1)$, dilatation (scale transformation), and special conformal transformation (SCT) [22, 23]. Each CFT operator must transform under irreducible representations of rotation and dilatation. The representations are labelled by the $\mathrm{SO}(d)$ spin $l$ and scaling dimension $\Delta$, respectively. A special kind of operators that are invariant under SCT, called ‘primaries’, deserve particular attention. By acting spatial derivatives on the primaries, their ‘descendants’ are obtained. Each operator in CFT is a linear combination of primaries and descendants. The conformal symmetry is the maximal spacetime symmetry (except supersymmetry) that a field theory can have. It gives powerful constraints on the property of the field theory. In particular, conformal symmetry uniquely determines the form of two-point (2-pt) and three-point (3-pt) correlation functions. The 3-pt correlator of three primary operators $\Phi_{i},\Phi_{j},\Phi_{k}$ contains a universal coefficient called the OPE coefficient $f_{\Phi_{i}\Phi_{j}\Phi_{k}}$. The collection of scaling dimensions and the OPE coefficients of primaries $\{\Delta_{\Phi_{i}},f_{\Phi_{i}\Phi_{j}\Phi_{k}}\}$ is called the conformal data. Theoretically, with full knowledge of the CFT data, an arbitrary correlation function of a CFT can be obtained [^2] The full knowledge is not often possible in practice, as the number of primaries is often infinite..
CFT has provided important insights into various aspects of theoretical physics. In condensed matter physics, it has produced useful predictions about the critical phenomena [24, 25, 26]. Many classical and quantum phase transitions are conjectured to have emergent conformal symmetry in the infrared (IR), i.e., at long wavelength or low energy. The universal critical exponents are directly determined by the scaling dimensions of the primary operators. E.g., in the 3d Ising transition that spontaneously breaks $\mathbb{Z}_{2}$ symmetry, most critical exponents are given by the scaling dimensions of the lowest $\mathbb{Z}_{2}$-odd operator $\sigma$ and $\mathbb{Z}_{2}$-even operator $\epsilon$, such as
@@
CFT is also closely related to string theory and quantum gravity in high-energy physics. In the string theory, CFT describes the 2d worldsheet [27] ; in quantum gravity, there is a conjectured duality between the gravity theory in $(d+1)$-dimensional anti-de Sitter (AdS) space in the bulk and a $d$-dimensional CFT on the boundary [28]. Moreover, CFT plays an important role in our understanding of quantum field theories. It describes many fixed points in the RG flow, and many QFTs can be seen as CFTs with perturbations. It also helps us understand how physics changes under a change of scale and reveals some fundamental structures of the RG flow [29].
In 2d CFTs, besides the global conformal symmetry $\mathrm{SO}(3,1)$, there also exists an infinite-dimensional local conformal symmetry [30, 31]. Altogether, they form the Virasoro algebra. The infinite-dimensional conformal algebra has made many theories exactly solvable, especially the rational theories such as the minimal models [32] and, more generally, the Wess-Zumino-Witten (WZW) theories [33, 34]. However, going to the higher dimensions, the CFTs are much less well-studied due to a much smaller conformal group. The existing methods include numerical conformal bootstrap and Monte Carlo lattice simulations. Numerical bootstrap bounds the conformal data by making use of consistency conditions such as reflection positivity and crossing symmetry, together with some information of the CFT such as the global symmetry and a certain amount of assumptions [35, 36]. It has achieved great success in 3d Ising [37, 38], $\mathrm{O}(N)$ Wilson-Fisher [39, 38], Gross-Neveu-Yukawa CFTs [40], etc. On the other hand, one can study a CFT by constructing a lattice model that goes through a phase transition in the corresponding universality class, and study the phase transition by Monte Carlo simulation. This method has achieved success in many phase transitions assuming conformal symmetry, e.g., the 3d Ising model [41]. However, the extraction of universal data usually involves complicated and expensive finite-size scaling [25, 42, 43], and only the lowest few CFT operators can be accessed.
Among these higher dimensional CFTs, we mainly focus on $d=3$, as many Lagrangians in $d\geq 4$ flow to free theories [26].
@@ 2.2 Fuzzy sphere
In addition to these existing approaches, the ‘fuzzy sphere regularisation’ has recently emerged as a new powerful method to study 3d CFTs. It involves studying interacting electrons moving on a sphere under the influence of a magnetic monopole at its centre.
The idea begins with putting an interacting quantum Hamiltonian on a 2-sphere $S^{2}$. This geometry preserves the full rotation symmetry (on the contrary, lattice models often only preserve a discrete subgroup). Moreover, when the system is tuned to a critical point or critical phase, combined with the time evolution direction, the system is described by a quantum field theory living on a generalised cylinder $S^{2}\times\mathbb{R}$, a manifold that is conformally equivalent to flat spacetime through the Weyl transformation
$(\hat{\mathbf{n}},\tau)\in S^{2}\times\mathbb{R}\ \longmapsto\ r\hat{\mathbf{n}}\in\mathbb{R}^{3},\qquad r=e^{\tau/R}$ (2)
where $R$ is the radius of the sphere. This conformal transformation maps each time slice of the cylinder to a concentric sphere in the flat spacetime.
Thanks to the conformal flatness that is not owned by other manifolds (e.g., a lattice model with periodic boundary condition lives on the torus $T^{2}$, which is not conformally flat), we can make use of some nice properties of conformal field theories. The most important one is the state-operator correspondance [44, 22, 23]. Specifically, there is a one-to-one correspondence between the eigenstates of the critical Hamiltonian on the sphere and the CFT operators. One can colloquially understand the state $|\Phi\rangle$ as the insertion of the corresponding operator $\Phi(0)$ at the origin point into the vacuum $|0\rangle$ : $|\Phi\rangle=\hat{\Phi}(0)|0\rangle$. The state and its corresponding operator have the same $\mathrm{SO}(3)$ spin and representation under global symmetry. More importantly, as the Weyl transformation maps the Hamiltonian $H$ that generates the time translation on the cylinder to the dilatation $D$ on the flat spacetime, the excitation energy of a state $|\Phi\rangle$ is proportional to the scaling dimension of the corresponding operator $\Delta_{\Phi}$
@@
Although the quantum Hamiltonians on a sphere enjoy the full rotation symmetry and the property of state-operator correspondence, it is difficult to put a lattice on the sphere due to the curvature (in particular the non-zero Euler characteristic), especially to recover an $\mathrm{SO}(3)$-symmetric thermodynamic limit [45]. An alternative way is to fuzzify the sphere [46]. We consider charged free particles moving on a sphere with a magnetic monopole with a flux $4\pi s$ ($s\in\mathbb{Z}/2$) placed at its centre. The monopole exerts a uniform magnetic field on the sphere, which modifies the single-particle Hamiltonian and the single-particle eigenstates. Now, the single particle eigenstates form highly degenerate spherical Landau levels [47, 48, 49, 50]. The lowest Landau level has a degeneracy $(2s+1)$. By setting the single-particle gap to be the leading energy scale, and projecting onto the lowest Landau level, we obtain a finite Hilbert space. For the purpose of numerical simulation, the system is analogous to a length-$(2s+1)$ chain with long-range interaction, where different Landau level orbitals behave like the lattice sites. The difference is that the $(2s+1)$ orbital forms a spin-$s$ representation of the $\mathrm{SO}(3)$ rotation group, and in this way, the continuous rotation symmetry is preserved. By putting multiple flavours on an orbital and adding interactions, various 3d CFTs can be realised. The interaction Hamiltonians are designed through matching the global symmetry and phase diagram. The word ‘fuzzy’ means the non-commutativity, in our case, due to the presence of magnetic field [46, 50]. The non-commutativity provides a natural length scale which serves as a UV regulator of the quantum field theory. The radius of the sphere scales as $R\sim\sqrt{s}$. The thermodynamic limit can be taken as $s\to\infty$, and we then recover a regular sphere without non-commutativity.
@@
The power of this approach has been first demonstrated in the context of the 3d Ising transition [1], where the presence of emergent conformal symmetry has been convincingly established, and a wealth of conformal data has been accurately computed. The study has then been extended to accessing various conformal data such as the OPE coefficients [2], correlation functions [3], entropic $F$-function [10], conformal generators [14, 15], studying conformal defects and boundaries such as the magnetic line defect [6, 9], various conformal boundaries in the 3d Ising CFT [12, 13], and realising various 3d CFTs such as Wilson-Fisher CFTs [8], $\mathrm{SO}(5)$ deconfined criticality [4], etc. In the following sections, we shall review the existing works, technical details and numerical methods.
@@ 3 Review of existing works
This seminal work opens a new avenue for studying 3d conformal field theories. After that, most of the research on the fuzzy sphere can roughly be categorised into three directions (with several exceptions) :
@@
    Realising various 3d CFTs, and
@@ 3.1 Accessing various conformal data
The first direction is to develop methods to calculate various universal data of 3d CFTs on the fuzzy sphere. Typically, these methods are tested with the simplest example of the 3d Ising CFT. For many of those CFT data, the fuzzy sphere is the first non-perturbative method to access them ; for the others, the fuzzy sphere has achieved great consistency with previous methods such as quantum Monte Carlo and conformal bootstrap. So far, the accessible CFT data include operator spectrum, OPE coefficients [2], correlation functions [3], entropic $F$-function [10] and conformal generators [14, 15].
@@
Apart from the operator spectrum, a wealth of CFT data can be obtained from the local operators. This work studies the local observables on the fuzzy sphere, including the density operators and certain four-fermion operators. These observables can be expressed as the linear combination of local scaling operators in the CFT. After a finite-size scaling with the data from different system sizes, the subleading contribution can be subtracted, and only the leading contribution is left. In this way, the lowest primaries in Ising CFT in each symmetry sector, viz. the $\mathbb{Z}_{2}$-odd $\sigma$ and the $\mathbb{Z}_{2}$-even $\epsilon$, can be realised. The OPE coefficients are then evaluated by taking the inner product of a fuzzy sphere local observable with two CFT states $\langle\Phi_{1}|\Phi_{2}(\hat{\mathbf{n}})|\Phi_{3}\rangle$. Hu et al. compute 17 OPE coefficients of low-lying CFT primary fields with high accuracy, including four that have not been reported before. The rest are consistent with numerical bootstrap results. It is also worth noting that this work starts to apply DMRG to the fuzzy sphere. The maximal system size is increased from $N_{m}=18$ by ED to $N_{m}=48$ by DMRG. The detail for calculating the OPE coefficient is presented in Section 4.6.
@@
In addition to the OPE coefficients, the local observables can also be used to calculate correlation functions. By taking the inner product of two local observables (density operators) at a time displacement $\langle\Phi_{1}|\Phi_{2}(\hat{\mathbf{n}}_{0})\Phi_{3}(\hat{\mathbf{n}},\tau)|\Phi_{4}\rangle$ with two CFT states, a general 4-pt function can be calculated. In practice, this piece of CFT data cannot be derived from the scaling dimensions and the OPEs due to the existence of infinitely many primaries. Han et al. calculate the 4-pt functions in the 3d Ising CFT with DMRG. A non-trivial check of conformality, the crossing symmetry, is verified for the correlator $\langle\sigma\sigma\sigma\sigma\rangle$. The special case — 2-pt functions by taking $\Phi_{1}=\Phi_{4}=\mathbb{I}$ — is also studied and compared with the expected results by conformal symmetry. The detail for calculating the correlation functions is presented in Section 4.6.
@@
where $\delta$ is a UV-regulator. The constant piece is known as the $F$-function of a 3d CFT. The $F$-function is proved to be RG-monotonic, i.e., along a renormalisation group flow from UV to IR, the value of $F$-function is non-increasing, analogous to the central charge in 2d CFTs. Despite its importance, it has never been calculated before through non-perturbative approaches in interacting 3d CFTs. This work has performed the first non-perturbative computation of $F$ function for the 3d Ising CFT on the fuzzy sphere. The sphere is cut in the real space into two crowns along a latitude circle $\theta$, and the entanglement entropy $S_{A}(\theta)$ as a function of $\theta$ is calculated [58, 59, 60, 61]. The $F$-function is extracted from the $S_{A}(\theta)$ in the vicinity of the equator, and the result yields $F=0.0612(5)$ after a finite-size scaling.
@@
[…]

### v3

@@ 1 Purpose and Outline
Since its proposal, the fuzzy sphere regularisation has made significant contributions to the study of 3D CFTs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. The Julia package FuzzifiED aims at simplifying the numerical calculations on the fuzzy sphere. It supports exact diagonalisation (ED) calculations, as well as the density matrix renormalisation group (DMRG) using the ITensor [35] library. FuzzifiED can also apply to generic fermionic and bosonic models. This package offers the following features :
@@
    Flexibility : FuzzifiED can reproduce nearly all the ED and DMRG results from fuzzy sphere research. Its flexible design also makes it straightforward to adapt to new models.
@@
  • In Section 4, we review the set-up of fuzzy sphere, the construction of interaction models and the extraction of CFT data.
@@
  • In Appendix B, we present the construction of the Landau level and pseudo-potentials on the torus.
@@
Part I Review of the Fuzzy Sphere
@@ 2.1 Conformal Field Theory
Conformal field theory (CFT) is one of the central topics of modern physics. It refers to a field theory that is invariant under conformal transformations that preserve the angles between vectors. In space-time dimension $d>2$, the global conformal symmetry transformations form a group $\mathrm{SO}(d+1,1)$, generated by translation, $\mathrm{SO}(d)$ rotation, [^1] Here we work in Euclidean signature. In Lorentzian signature it is the Lorentz transformation $\mathrm{SO}(1,d-1)$ dilatation (scale transformation), and special conformal transformation (SCT) [37, 38]. Each CFT operator must transform under irreducible representations of rotation and dilatation. The representations are labelled by the $\mathrm{SO}(d)$ spin $l$ and scaling dimension $\Delta$, respectively. A special kind of operators that are invariant under SCT, called “primaries,” deserve particular attention. By acting spatial derivatives on the primaries, their “descendants” are obtained. Each local operator in a CFT is a linear combination of primaries and descendants. The conformal symmetry is the maximal space-time symmetry (except supersymmetry) that a field theory can have. It gives powerful constraints on the property of the field theory. In particular, conformal symmetry uniquely determines the form of two-point (2-pt) and three-point (3-pt) correlation functions. The 3-pt correlator of three primary operators $\Phi_{i},\Phi_{j},\Phi_{k}$ contains a universal coefficient called the OPE coefficient $f_{\Phi_{i}\Phi_{j}\Phi_{k}}$. The collection of the scaling dimensions and the OPE coefficients of primaries $\{\Delta_{\Phi_{i}},f_{\Phi_{i}\Phi_{j}\Phi_{k}}\}$ is called the conformal data. Theoretically, with full knowledge of the CFT data, an arbitrary correlation function of a CFT can be obtained. [^2] The full knowledge is not often possible in practice, as the number of primaries is often infinite.
CFT has provided important insights into various aspects of theoretical physics. It has produced useful predictions about the critical phenomena [39, 40, 41] in condensed-matter physics. Many classical and quantum phase transitions are conjectured to have emergent conformal symmetry in the infra-red (IR), i. e. at long wave-length or low energy. The universal critical exponents are directly determined by the scaling dimensions of the primary operators. E. g., in the 3D Ising transition that spontaneously breaks a $\mathbb{Z}_{2}$ symmetry, most critical exponents are given by the scaling dimensions of the lowest $\mathbb{Z}_{2}$-odd operator $\sigma$ and $\mathbb{Z}_{2}$-even operator $\epsilon$, such as
@@
CFT is also closely related to string theory and quantum gravity in high-energy physics. In the string theory, CFT describes the 2D worldsheet [42] ; in quantum gravity, there is a conjectured duality between the gravity theory in $(d+1)$-dimensional anti-de Sitter (AdS) space in the bulk and a $d$-dimensional CFT on the boundary [43]. Moreover, CFT plays an important role in understanding of quantum field theories in general. It describes many fixed points in the renormalisation-group (RG) flow, and many QFTs can be seen as CFTs with perturbations. It also helps us understand how physics changes under a change of scale and reveals some fundamental structures of the RG flow [44].
In 2D CFTs, besides the global conformal symmetry $\mathrm{SO}(3,1)$, there also exists an infinite-dimensional local conformal symmetry [45, 46]. Altogether, they form the Virasoro algebra. The infinite-dimensional conformal algebra has made many theories exactly solvable, especially the rational theories such as the minimal models [47] and, more generally, the Wess-Zumino-Witten (WZW) theories [48, 49]. However, going to higher dimensions, the CFTs are less well-studied due to a much smaller conformal group. The existing methods include numerical conformal bootstrap and Monte Carlo lattice simulations. Numerical bootstrap bounds the conformal data by making use of consistency conditions such as reflection positivity and crossing symmetry, together with some information of the CFT such as the global symmetry and a certain amount of assumptions [50, 51]. It has achieved great success in 3D Ising [52, 53], $\mathrm{O}(N)$ Wilson-Fisher [54, 55], Gross-Neveu-Yukawa CFTs [56], etc. On the other hand, one can study a CFT by constructing a lattice model that goes through a phase transition in the corresponding universality class, and study the phase transition by Monte Carlo simulation. This method has achieved success in many phase transitions assuming conformal symmetry, e. g. the 3D Ising model [57]. However, the extraction of universal data usually involves complicated and computation-intensive finite-size scaling [40, 58, 59], and only the lowest few CFT operators can be accessed.
Among these higher dimensional CFTs, we mainly focus on three space-time dimensions.
@@ 2.2 Fuzzy Sphere
In addition to these existing approaches, the “fuzzy sphere regularisation” has recently emerged as a new powerful method to study 3D CFTs. It involves studying interacting electrons moving on a sphere under the influence of a magnetic monopole at its centre.
The idea begins with putting an interacting quantum Hamiltonian on a 2-sphere $S^{2}$. This geometry preserves the full rotation symmetry (on the contrary, lattice models often only preserve a discrete subgroup). Moreover, when the system is tuned to a critical point or critical phase, combined with the time evolution direction, the system is described by a quantum field theory living on a generalised cylinder $S^{2}\times\mathbb{R}$, a manifold that is conformally equivalent to flat space-time through the Weyl transformation
$(\mathbf{r},\tau)\in S^{2}\times\mathbb{R}\ \longmapsto\ e^{\tau/R}\hat{\mathbf{n}}\in\mathbb{R}^{3},$ (2)
where $R$ is the radius of the sphere and $\hat{\mathbf{n}}$ is the unit vector corresponding to $\mathbf{r}$. This conformal transformation maps each time slice of the cylinder to a co-centric sphere in the flat space-time.
Thanks to the conformal flatness that is not owned by other manifolds (e. g., a lattice model with periodic boundary condition lives on the torus $T^{2}$, which is not conformally flat), we can make use of some nice properties of conformal field theories. The most important one is the state-operator correspondance [60, 37, 38]. Specifically, there is a one-to-one correspondence between the eigenstates of the critical Hamiltonian on the sphere and the CFT operators. One can colloquially understand the state $|\Phi\rangle$ as the insertion of the corresponding operator $\Phi(0)$ at the origin point into the vacuum $|0\rangle$ : $|\Phi\rangle=\hat{\Phi}(0)|0\rangle$. The state and its corresponding operator have the same $\mathrm{SO}(3)$ spin and representation under global symmetry. More importantly, as the Weyl transformation maps the Hamiltonian $H$ that generates the time translation on the cylinder to the dilatation $D$ on the flat space-time, the excitation energy of a state $|\Phi\rangle$ is proportional to the scaling dimension of the corresponding operator $\Delta_{\Phi}$
@@
Although the quantum Hamiltonians on a sphere enjoy the full rotation symmetry and the property of state-operator correspondence, it is difficult to put a lattice on the sphere due to the curvature (in particular the non-zero Euler characteristic), especially to recover an $\mathrm{SO}(3)$-symmetric thermodynamic limit [61]. An alternative way is to fuzzify the sphere [62]. We consider charged free particles moving on a sphere with a magnetic monopole with a flux $4\pi s$ ($s\in\mathbb{Z}/2$) placed at its centre. The monopole exerts a uniform magnetic field on the sphere, which modifies the single-particle Hamiltonian and the single-particle eigenstates. Now, the single particle eigenstates form highly degenerate spherical Landau levels [63, 64, 65, 66]. The lowest Landau level has a degeneracy $(2s+1)$. By setting the single-particle gap to be the leading energy scale, and projecting onto the lowest Landau level, we obtain a finite Hilbert space. For the purpose of numerical simulation, the system is analogous to a length-$(2s+1)$ chain with long-range interaction, where different Landau level orbitals behave like the lattice sites. The difference is that the $(2s+1)$ orbitals form a spin-$s$ representation of the $\mathrm{SO}(3)$ rotation group, and in this way, the continuous rotation symmetry is preserved. By putting multiple flavours on an orbital and adding interactions, various 3D CFTs can be realised. The interaction Hamiltonians are designed through matching the global symmetry and phase diagram. The word “fuzzy” means the non-commutativity, in our case, due to the presence of magnetic field [62, 66]. The non-commutativity provides a natural length scale which serves as a UV regulator of the quantum field theory. The radius of the sphere scales as $R\sim\sqrt{s}$. The thermodynamic limit can be taken as $s\to\infty$, and we then recover a regular sphere without non-commutativity.
@@
The power of this approach has been first demonstrated in the context of the 3D Ising transition [1], where the presence of emergent conformal symmetry has been convincingly established, and a wealth of conformal data has been accurately computed. The study has then been extended to accessing various conformal data such as the OPE coefficients [2], correlation functions [3], entropic $F$-function [10], conformal generators [14, 15] and the cross-cap coefficients [28], developing techniques to improve the numerical precision, such as quantum Monte Carlo [7], conformal perturbation [20] and the finite-size scaling of the ground-state energy [31], realising various 3D CFTs such as the free scalar [24, 25], Wilson-Fisher CFTs [8, 32, 33], $\mathrm{SO}(5)$ [4] and $\mathrm{O}(4)$ [26] deconfined criticality, $\mathrm{Sp}(N)$-symmetric CFTs [16], multi-flavour $\mathrm{SU}(2)$ QCD [34], 3-state Potts model [18] and Yang-Lee non-unitary CFT [21, 22, 23], exploring fractional quantum Hall transitions, such as the Ising CFT on the FQHE states [17], the confinement transition of the $\nu=1/2$ bosonic Laughlin state [27], the free Marjorana fermion theory [29], and the transition between bosonic Pfaffian and the Halperin 220 state [30], studying conformal defects and boundaries such as the pinning-field defect [6, 9], various conformal boundaries in the 3D Ising CFT [12, 13], and lower-dimensional CFTs like 2D Ising on the fuzzy circle [19]. In the following sections, we shall review the existing works, technical details and numerical methods.
@@ 3 Review of Existing Works
This seminal work opens a new avenue for studying 3D conformal field theories. After that, most of the research on the fuzzy sphere can roughly be categorised into the following directions :
@@
    Developing techniques to improve the numerical precision,
@@
    Realising various 3D CFTs,
  4. 4.
    Exploring fractional quantum Hall transitions and
  5. 5.
@@ 3.1 Accessing Various Conformal Data
The first direction is to develop methods to calculate various universal data of 3D CFTs on the fuzzy sphere. Typically, these methods are tested with the simplest example of the 3D Ising CFT. For many of those CFT data, the fuzzy sphere is the first non-perturbative method to access them ; for the others, the fuzzy sphere has achieved great consistency with previous methods such as quantum Monte Carlo and conformal bootstrap. So far, the accessible CFT data include operator spectrum, OPE coefficients [2], correlation functions [3], entropic $F$-function [10], conformal generators [14, 15], the cross-cap coefficients [28].
@@
Apart from the operator spectrum, a wealth of CFT data can be obtained from the local operators. This work studies the local observables on the fuzzy sphere, including the density operators and certain four-fermion operators. These observables can be expressed as linear combinations of local scaling operators in the CFT. After a finite-size scaling with the data from different system sizes, the sub-leading contribution can be subtracted, and only the leading contribution is left. In this way, the lowest primaries in the Ising CFT in each symmetry sector, viz. the $\mathbb{Z}_{2}$-odd $\sigma$ and the $\mathbb{Z}_{2}$-even $\epsilon$, can be realised. The OPE coefficients are then evaluated by taking the inner product of a fuzzy sphere local observable with two CFT states $\langle\Phi_{1}|\Phi_{2}(\mathbf{r})|\Phi_{3}\rangle$. Hu et al. compute 17 OPE coefficients of low-lying CFT primary fields with high accuracy, including four that have not been reported before. The rest are consistent with numerical bootstrap results. It is also worth noting that this work starts to apply DMRG to the fuzzy sphere. The maximal system size is increased from $N_{m}=18$ by ED to $N_{m}=48$ by DMRG. The detail for calculating the OPE coefficient is presented in Section 4.6.
@@
In addition to the OPE coefficients, the local observables can also be used to calculate correlation functions. By taking the inner product of two local observables (density operators) at a time displacement $\langle\Phi_{1}|\Phi_{2}(\mathbf{r}_{0})\Phi_{3}(\mathbf{r},\tau)|\Phi_{4}\rangle$ with two CFT states, a general 4-pt function can be calculated. In practice, this piece of CFT data cannot be derived from the scaling dimensions and the OPEs due to the existence of infinitely many primaries. Han et al. calculate the 4-pt functions in the 3D Ising CFT with DMRG. A non-trivial check of conformality, the crossing symmetry, is verified for the correlator $\langle\sigma\sigma\sigma\sigma\rangle$. The special case — 2-pt functions by taking $\Phi_{1}=\Phi_{4}=\mathbb{I}$ — is also studied and compared with the expected results by conformal symmetry. The detail for calculating the correlation functions is presented in Section 4.6.
@@
where $\delta$ is a UV-regulator. The constant piece is known as the $F$-function of a 3D CFT. The $F$-function is proved to be RG-monotonic, i. e., along a RG flow from UV to IR, the value of $F$-function is non-increasing, analogous to the central charge in 2D CFTs. Despite its importance, it has never been calculated before through non-perturbative approaches in interacting 3D CFTs. This work has performed the first non-perturbative computation of $F$ function for the 3D Ising CFT on the fuzzy sphere. The sphere is cut in the real space into two crowns along a latitude circle $\theta$, and the entanglement entropy $S_{A}(\theta)$ as a function of $\theta$ is calculated [74, 75, 76, 77]. The $F$-function is extracted from the $S_{A}(\theta)$ in the vicinity of the equator, and the result yields $F=0.0612(5)$ after a finite-size scaling.
@@
Within the generators of conformal symmetry, the $\mathrm{SO}(3)$ rotation and the dilatation are manifest and act as rotation and time translation on the fuzzy sphere. The rest two, viz. translation $P^{\mu}$ and special conformal transformation (SCT) $K^{\mu}$, need to be emergent in the IR at the conformal point. It is worthwhile to construct these IR generators by the UV operators on the fuzzy sphere. These works investigate such construction with the help of stress tensor $T^{\mu\nu}$. The time component $T^{\tau\tau}$ of stress tensor equals the Hamiltonian density $\mathscr{H}$ and it integrates into the generator $\Lambda^{\mu}=P^{\mu}+K^{\mu}=\int\mathrm{d}^{2}\mathbf{r}\,2n^{\mu}\mathscr{H}$. The action of this generator sends a scaling operator to other operators in the same multiplet, with the number of spatial derivatives increased or decreased by one. Fardelli et al. and Fan calculate the matrix elements of the generators $\Lambda^{\mu}$ and find good agreement with the theoretical values in the CFT, which is another non-trivial verification of conformal symmetry. Furthermore, the separate generators $P^{\mu}$ and $K^{\mu}$ can be obtained by considering the commutator $H,\Lambda^{\mu}]$, which is useful in determining the primaries. The detail for constructing the conformal generators is presented in Section 4.7.
The cross-cap coefficients [28]
This work focuses on the 3D CFTs on the space-time manifold of the the real projective space $\mathbb{R}\mathrm{P}^{3}$ obtained by identifying antipodal points on $S^{2}$. On $\mathbb{R}\mathrm{P}^{3}$, one-point functions of scalar primary fields are generally non-vanishing and encodes the “cross-cap coefficients.” Dong et al. extracts the cross-cap coefficients of the 3D Ising CFT through simulating microscopic models on the lattice models on polyhedrons and continuum models in Landau levels and entangle the degrees of freedom at anti-podal points in Bell-type states.
3.2 Developing Techniques to Improve the Numerical Precision
There have also been techniques to improve the precision through accessing larger system size and detailed analysis of finite-size data, such as quantum Monte Carlo simulation [7], conformal perturbation [20], and finite-size scaling the ground-state energy [31].
Quantum Monte Carlo on the fuzzy sphere [7]
Until this work, the numerical methods that have been applied to the fuzzy sphere include exact diagonalisation (ED) and density matrix renormalisation group (DMRG). Hofmann et al. further present the numerical studies of the fuzzy sphere with quantum Monte Carlo (QMC) simulation, which is known for its potential for studying criticalities in $(2+1)$ dimensions at larger system size. Specifically, they make use of the determinant quantum Monte Carlo (DQMC) method that converts the simulation of fermions into the simulation of bosonic auxiliary fields. To overcome the sign problem, they consider two copies of the original model and construct the Ising CFT on a 4-flavour model. They determine the lowest energy spectra within each symmetry sector by calculating the imaginary-time-displaced correlation functions. They also calculate the equal-time correlation functions and compare them with the 2-pt functions of CFT.
Conformal perturbation [20]
The energy spectrum calculated numerically at a finite size does not coincide with that of the CFT. Part of the finite-size correction comes from the higher irrelevant operators that are not precisely tuned to zero (e. g., in the Ising CFT, the irrelevant operators include $\epsilon^{\prime},C_{\mu\nu\rho\sigma},T^{\prime}_{\mu\nu}$, etc., and the lowest singlets $\epsilon$ and $\epsilon^{\prime}$ are tuned away through the two parameters). These irrelevant operators exert perturbations on the states and their energies. This work captures this kind of correction using the conformal perturbation theory. By making use of the fact that the corrections from an irrelevant operator on the energy of the primary and its descendants are not independent, the coefficients of the irrelevant operators can be fitted and their corrections can be removed.
The conformal perturbation theory is first studied on the 2D CFTs [78, 79] and the icosahedron [5] and then applied to the fuzzy sphere. This work opens up a new route to improving the precision of scaling dimensions on the fuzzy sphere by making better use of the existing data. The method to partly remove the finite-size correction through conformal perturbation theory is widely used by the following works.
Finite-size scaling of the ground-state energy [31]
This work proposes a new approach to locate the phase-transition line from a finite-size scaling analysis of its ground-state energy with the example of the 3D Ising CFT. Wiese performs a finite-size scaling with the ansatz $E_{\text{GS}}=E_{0}R^{2}+E_{1}+E_{3/2}R^{-1}$ and identify the minima of $E_{3/2}/E_{0}$ as the critical curve and the “sweet spot.”